### v3

# GPSMamba: A Global Phase and Spectral Prompt-guided Mamba for Infrared Image Super-Resolution

Abstract
Infrared Image Super-Resolution (IRSR) is challenged by the low contrast and sparse textures of infrared data, requiring robust long-range modeling to maintain global coherence. While State-Space Models like Mamba offer proficiency in modeling long-range dependencies for this task, their inherent 1D causal scanning mechanism fragments the global context of 2D images, hindering fine-detail restoration. To address this, we propose Global Phase and Spectral Prompt-guided Mamba (GPSMamba), a framework that synergizes architectural guidance with non-causal supervision. First, our Adaptive Semantic-Frequency State Space Module (ASF-SSM) injects a fused semantic-frequency prompt directly into the Mamba block, integrating non-local context to guide reconstruction. Then, a novel Thermal-Spectral Attention and Phase Consistency Loss provides explicit, non-causal supervision to enforce global structural and spectral fidelity. By combining these two innovations, our work presents a systematic strategy to mitigate the limitations of causal modeling. Extensive experiments demonstrate that GPSMamba achieves state-of-the-art performance, validating our approach as a powerful new paradigm for infrared image restoration. Code is available at https://github.com/yongsongH/GPSMamba.
Introduction
Figure 1: An overview of our proposed GPSMamba for IRSR. (a) Classical Mamba(Xiao et al. 2025). The fixed, causal scan path of classical Mamba struggles to capture long-range dependencies in low-contrast infrared images, limiting reconstruction fidelity. (b) GPSMamba. To overcome this, GPSMamba introduces two key innovations: 1) The Adaptive Semantic Fusion State Space Model (ASF-SSM) injects a global frequency-domain prompt ($P_{\text{fused}}$) via an adaptive scan path, breaking the rigid causal chain to integrate non-local context. 2) A novel Thermal-Spectral Attention and Phase Consistency (TSAPC) Loss ($\mathcal{L}_{T}$) provides explicit, non-causal supervision to enforce global structural and spectral fidelity.
Infrared (IR) imaging is indispensable in applications such as security surveillance, remote sensing(Zhu et al. 2022), and autonomous driving(Tan et al. 2024), particularly in environments where visible-light imaging is compromised by poor illumination or adverse weather conditions(Fang et al. 2024). However, raw IR images frequently suffer from low spatial resolution, noise contamination, and poor contrast, stemming from sensor hardware limitations and atmospheric interference(Huang et al. 2023). The objective of IRSR is to reconstruct a high-resolution (HR) image, $I_{HR}\in\mathbb{R}^{H\times W\times C}$, from its low-resolution (LR) counterpart, $I_{LR}\in\mathbb{R}^{h\times w\times C}$, where $H,W,C$ denote the height, width, and channel count, and $s$ is the scaling factor ($h=H/s,w=W/s$). This process can be formulated as learning a mapping function $F$ such that $I_{SR}=F(I_{LR})$, where $I_{SR}$ is the reconstructed image that aims to approximate the $I_{HR}$.
Unlike natural image super-resolution, IRSR presents unique challenges due to the intrinsic properties of thermal radiation(Huang et al. 2025). IR images are often characterized by low contrast, homogeneous backgrounds, and sparse high-frequency details, making it difficult for conventional models to restore fine textures. Furthermore, targets of interest may be spatially distant from each other, demanding that the reconstruction model possesses a superior capability for long-range contextual awareness to ensure global structural coherence(Li et al. 2025). With the advent of deep learning, data-driven methods have become the dominant paradigm in super-resolution(Liu et al. 2022a). Their primary advantage lies in the ability to learn the complex, non-linear mapping from LR to HR feature spaces directly from data, obviating the need for meticulously designed, hand-crafted priors(Wang et al. 2020). This has established deep learning as the mainstream approach for image restoration tasks. Recently, State-Space Models (SSMs)(Guo et al. 2024; Zhu et al. 2024), particularly Mamba(Gu and Dao 2023), have emerged as a powerful architecture for sequence modeling. Their ability to capture long-range dependencies with linear computational complexity presents a compelling alternative to high-complexity Transformers and GANs, which can be prone to training instability. The potential of SSMs has been successfully demonstrated in image restoration. For instance, MambaIR(Guo et al. 2024) first validated the effectiveness of Mamba in this domain. MambaIRv2(Guo et al. 2025) further enhanced global context modeling by using similarity maps to guide the model’s attention toward semantically relevant regions. Most recently, IRSRMamba(Huang et al. 2025) achieved state-of-the-art (SOTA) performance on several IRSR benchmarks by integrating global priors from the frequency domain.
Despite this progress, we observe that the intrinsic causal modeling of SSMs still poses a fundamental challenge for IRSR. The fixed, 1D causal scanning mechanism of Mamba inherently flattens a 2D image, fragmenting its global structural information into discrete and non-adjacent patches(Guo et al. 2025; Xiao et al. 2025; Shao et al. 2025). To faithfully restore fine-grained details and maintain global coherence in IR images, it is crucial to develop an SSM-based method that can better aggregate information from semantically relevant regions across the entire image (see Fig.1). In this work, we propose GPSMamba to address this limitation. As illustrated in Fig. 1, our approach introduces two core innovations: 1) a Mamba architecture enhanced by a global frequency-domain prompt to guide the reconstruction with non-local context, and 2) a novel non-causal loss function designed to explicitly enforce global structural and spectral fidelity. More details will be shown in the method section. Our contributions are as follows:
  • An Adaptive Semantic-Frequency State Space Module that injects a global frequency-domain prompt to guide the Mamba block, breaking its causal modeling limitations to enhance non-local context fusion.
  • We also introduce a novel Thermal-Spectral Attention and Phase Consistency Loss ($\mathcal{L}_{T}$), which provides explicit, non-causal supervision. This loss directly compensates for the information fragmentation caused by causal scanning by enforcing global structural and spectral fidelity across the image.
  • GPSMamba achieves new SOTA performance on benchmark IRSR datasets, establishing a more effective paradigm for Mamba-based infrared image restoration.
Related Work
Figure 2: The overall architecture of our proposed GPSMamba model. (a) GPSMamba Body: The top-level architecture processes $I_{LR}$ through a series of Adaptive Semantic-Frequency State Space Blocks (ASF-SSB) to produce the $I_{SR}$. The entire network is trained end-to-end under the supervision of our proposed TSAPC Loss ($\mathcal{L}_{T}$). (b) Adaptive Semantic-Frequency State Space Block (ASF-SSB): This is the main building block of the network. It takes an input feature map $X$ and refines it to produce an output $Y$ via a residual connection. Each ASF-SSB contains several of our core ASF-SSM modules, which are responsible for the SSMs scan mechanism. (c) Axial Shifted Fusion State Space Module (ASF-SSM): It first dynamically generates a complete prompt tensor $P^{\prime}$ by fusing two components: (1) a content-aware spatial prompt, $P_{\text{spatial}}$, selected from a learnable prompt pool $P$ using a Gumbel-Softmax router, and (2) a global context prompt $P_{\text{global}}$ derived from frequency-domain features $F_{\text{freq}}$ via self-attention. Then, at each timestep $t$, a time-specific vector is selected, $P_{\text{fused}}(t)=P^{\prime}_{t}$, and injected to guide the state transition. This fusion allows the module to model both local semantics and global structure simultaneously.
IRSR is an inherently challenging task due to the unique properties of IR data, such as low contrast, high noise levels, and large homogeneous regions with sparse textural details. Deep learning approaches have primarily followed two main architectural paradigms to address this. While Convolutional Neural Network (CNN) based methods are efficient at extracting local features, their limited receptive fields hinder the modeling of long-range dependencies, which is crucial for maintaining structural coherence in IR images(Shi et al. 2024; Qin et al. 2024b; Wu et al. 2023). To overcome this limitation, Transformer-based architectures were introduced, leveraging self-attention mechanisms to effectively capture global context and achieve state-of-the-art performance. However, the quadratic computational complexity of self-attention makes these models impractical for HR applications and resource-constrained scenarios(Qin et al. 2024b). This establishes a critical trade-off between global modeling capability and computational efficiency. Therefore, there is a pressing need for a new architecture that can efficiently model long-range dependencies for high-fidelity IRSR, motivating our exploration of SSMs.
State-Space Models for Image Restoration
SSMs, particularly the recently proposed Mamba, have emerged as a compelling alternative, offering linear complexity while maintaining the ability to model long-range dependencies. Mamba is built upon a structured SSMs that map a 1D input sequence $x_{t}$ to an output $y(t)$ via a latent state $h_{t}$, governed by the state equations:
$h_{t}=\mathbf{A}h_{t-1}+\mathbf{B}x_{t},\quad y_{t}=\mathbf{C}h_{t}$ (1)
where its core innovation, the Selective Scan Mechanism (S6), makes the state matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ input-dependent(Zhang et al. 2024a; Wang et al. 2024). This allows the model to dynamically focus on relevant information within the sequence. The efficiency and power of Mamba have led to its rapid adoption in computer vision. For image restoration, models like MambaIR(Guo et al. 2024), MambaIRv2(Guo et al. 2025), and IRSRMamba(Huang et al. 2025) have successfully applied SSMs, achieving SOTA results. These methods typically adapt Mamba to 2D data by flattening image patches into 1D sequences and processing them with bidirectional scans. However, this mandatory flattening and the inherent causal nature of the scanning mechanism can disrupt the spatial contiguity of the 2D image structure(Xiao et al. 2025; Shao et al. 2025). This fragmentation of global context is a significant drawback, especially for IRSR, where restoring sparse textures relies heavily on a holistic understanding of the entire scene. While existing works have attempted to mitigate this with multi-scale feature fusion or semantic guidance, these are often post-hoc corrections. Our work, GPSMamba, addresses this fundamental issue by directly integrating a global prompt into the SSM’s scanning process, enabling non-local information to guide feature extraction from the outset.
Global Priors and Guidance for Restoration
Incorporating priors is a well-established strategy to regularize the ill-posed nature of image restoration. Frequency-domain information, for example, serves as a powerful global prior. The Fourier spectrum decouples an image’s global structure (magnitude) from its fine-grained details (phase), which has been exploited by various methods to improve reconstruction fidelity(Xiao et al. 2024; Xiong et al. 2024). Other forms of guidance have also been explored. For instance, some works have designed uncertainty-driven loss functions to make the training process focus on more challenging regions(Peng, Xie, and Bian 2024). More recently, techniques like all-in-one restoration models have been developed by adapting large pre-trained networks with minimal overhead(Qin et al. 2024a; Yao et al. 2024). Our work introduces a novel form of guidance by framing the global prior as a dynamic prompt. Unlike methods that use frequency information as a regularization loss or in a separate processing branch, we inject a frequency-domain prompt directly into the core of our proposed ASF-SSM. This prompt actively modulates the state transitions within the Mamba block, making the feature extraction process aware of the global image context at every step. This tight integration of prompt-based guidance with the SSM architecture is a key distinction from previous works.
Method
Our proposed method, GPSMamba, is a hierarchical network designed for IRSR, with its overall architecture illustrated in Fig. 2. The network first employs a shallow convolutional layer to extract low-level features from the $I_{LR}$ input. The core of the model is a deep feature extraction body composed of several stacked ASF-SSBs. The central innovation lies within our ASF-SSM, which enhances the standard state-space mechanism with a dynamic, content-aware prompt generated from both semantic and spectral information. Finally, a reconstruction module upscales the learned deep features to produce the $I_{SR}$. To guide the network’s training, we introduce a novel objective function, the Thermal-Spectral Attention and Phase Consistency (TSAPC) Loss ($\mathcal{L}_{T}$). The following subsections provide a detailed exposition of our proposed ASF-SSM and the formulation of the TSAPC Loss.
Algorithm 1 ASF-SSM Core Processing
1: $X_{\text{semantic}}\in\mathbb{R}^{B\times N\times C}$, $P\in\mathbb{R}^{B\times N\times T}$, $Y_{\text{scan}}$
2: $Y\in\mathbb{R}^{B\times N\times C}$
[…]
13:for $t=1\ldots N$ do
14:  $P_{\text{fused}}(t)\leftarrow P^{\prime}(t)$
15:  $C_{s}(t)\leftarrow C_{\text{raw}}(t)+P_{\text{fused}}(t)$
16:  $h_{t}\leftarrow A(t)\odot h_{t-1}+B(t)\odot X_{\text{semantic}}(t)$
17:  $y_{t}\leftarrow C_{s}\odot h_{t}$
18:  Append $y_{t}$ to $Y_{\text{scan}}$
[…]
20:return $Y$
Adaptive Semantic-Frequency State Space Module
The Adaptive Semantic-Frequency State Space Module (ASF-SSM) is the cornerstone of our architecture, designed to overcome a key limitation of standard SSMs. While models like Mamba achieve linear complexity by processing flattened image patches as 1D sequences, this serialization can disrupt the 2D spatial relationships and global frequency patterns crucial for image understanding. Our ASF-SSM addresses this by enriching the selective scan mechanism with a dynamic, multi-domain prompt that guides state transitions based on both local semantics and global context. The ASF-SSM, detailed in Algorithm 1, comprises two synergistic stages: (a) Fused Prompt Generation, and (b) Prompt-Guided State Space Model.
Fused Prompt Generation. We generate a comprehensive prompt tensor $P^{\prime}$ by fusing local and global context. This process begins with the input feature sequence $X_{\text{semantic}}\in\mathbb{R}^{B\times N\times C}$, where $B,N,$ and $C$ are the batch, sequence length, and channel dimensions, respectively. To obtain the local component, a content-aware spatial prompt $P_{\text{spatial}}$, we first employ an MLP to predict routing logits from $X_{\text{semantic}}$. These logits are converted into a discrete, one-hot routing matrix $L\in\mathbb{R}^{B\times N\times T}$ via the Gumbel-Softmax function, where $T$ is the size of a learnable prompt pool $P$. The spatial prompt is then computed as $P_{\text{spatial}}=L\cdot P$. Concurrently, to integrate global context, the input $X_{\text{semantic}}$ is first reshaped and then transformed into the frequency domain via a Fast Fourier Transform (FFT), yielding frequency features $F_{\text{freq}}$. Inspired by the ability of the frequency domain to capture holistic image properties, we process these features through an attention mechanism. Specifically, $F_{\text{freq}}$ is linearly projected into query ($Q$), key ($K$), and value ($V$) tensors to compute a context prompt, $P_{\text{global}}$. These two prompts, computed in parallel, are then fused via element-wise addition to create the final, comprehensive prompt tensor: $P^{\prime}\leftarrow P_{\text{spatial}}+P_{\text{global}}$.
Prompt-Guided State Space Model. Our core innovation is to guide the SSM’s selective scan by dynamically modulating its output projection matrix $C_{s}(t)$. We first compute a comprehensive prompt tensor $P^{\prime}$ by fusing spatial and global context prompts, $P^{\prime}=P_{\text{spatial}}+P_{\text{global}}$.
To break the rigid causal chain, we introduce a semantic-guided scan. During this scan, the time-specific vector $P_{\text{fused}}(t)$ is selected from the tensor $P^{\prime}$ according to this semantic order, which is determined by sorting tokens based on the argmax of the routing matrix $L$ (Guo et al. 2025). This prompt is then injected into the output projection:
$C_{s}(t)=C_{s}^{\mathrm{raw}}(t)+P_{\text{fused}}(t)$ (2)
The other dynamic SSM parameters ($\Delta$, $B(t)$, $C_{s}^{\mathrm{raw}}(t)$) are derived from a linear projection of the input sequence, $X_{\text{in}}=\text{Linear}(X_{\text{semantic}})$, followed by a split operation. The state update and output generation then follow the standard discretized SSM formulation, with the input sequence $X_{\text{semantic}}$ along the same semantic path:
$h_{t}$ $=A(t)\odot h_{t-1}+B(t)\odot X_{\text{semantic}}(t)$ (3)
$y_{t}$ $=C_{s}(t)\odot h_{t}$ (4)
The resulting output sequence $Y_{\text{scan}}=[y_{1},\ldots,y_{N}]$ is then passed through a final LayerNorm and linear projection, before being reordered back to its original spatial arrangement to produce the module’s output $Y$.
[…]
Scale Methods # Params. (K) result-A result-C CVC10
[…]
$\times 2$ EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 39.0493 11.8196 0.9414 39.8902 8.9865 0.9528 44.1770 2.7845 0.9713
ESRGAN[ECCVW 2018] (Wang et al. 2018) 16,661 38.7738 12.5212 0.9384 39.6111 9.5793 0.9500 44.0974 2.8477 0.9709
FSRCNN[ECCV 2016] (Dong et al. 2016) 475 39.1175 11.3761 0.9426 39.9858 8.6899 0.9535 44.1253 2.8162 0.9710
SRGAN[CVPR 2017] (Ledig et al. 2017) 1,370 39.0401 11.9024 0.9414 39.8678 9.0586 0.9527 44.1736 2.7851 0.9713
SwinIR[ICCV 2021] (Liang et al. 2021) 11,752 38.6899 12.5694 0.9374 39.5215 9.6530 0.9492 43.9980 2.8926 0.9704
SRCNN[T-PAMI 2015] (Dong et al. 2015) 57 38.9671 11.7216 0.9414 39.8642 8.8857 0.9524 44.0038 2.9084 0.9707
RCAN[ECCV 2018] (Zhang et al. 2018) 12,467 38.8145 12.4926 0.9391 39.7075 9.4220 0.9511 44.1205 2.8170 0.9713
PSRGAN[SPL 2021] (Huang et al. 2021) 2,414 39.2146 11.2409 0.9429 40.0543 8.6101 0.9539 44.2377 2.7454 0.9713
ShuffleMixer(tiny)[NIPS’22] (Sun et al. 2022)) 108 39.0465 11.7605 0.9414 39.8766 8.9680 0.9527 44.1408 2.8113 0.9713
ShuffleMixer (base)[NIPS’22] (Sun et al. 2022) 121 38.8066 12.3718 0.9388 39.6347 9.4864 0.9503 44.0357 2.8809 0.9710
HAT[CVPR 2023] (Chen et al. 2023) 20,624 38.7754 12.4528 0.9384 39.6346 9.5132 0.9500 44.1080 2.8244 0.9709
RGT[ICLR 2024] (Chen et al. 2024) 10,051 39.1642 11.3382 0.9429 40.0522 8.6033 0.9540 44.2311 2.7358 0.9717
MambaIR [ECCV 2024 SOTA] (Guo et al. 2024) 20,421 39.1761 11.2081 0.9437 40.1399 8.4798 0.9544 44.4181 2.6076 0.9720
ATD[CVPR 2024] (Zhang et al. 2024b) 753 39.0453 11.5702 0.9432 40.0375 8.6155 0.9542 44.1901 2.7737 0.9711
CATANet [CVPR 2025 SOTA] (Liu et al. 2025) 477 39.0886 11.4332 0.9430 40.0064 8.6387 0.9539 44.1656 2.7858 0.9712
MambaOut[CVPR 2025 SOTA](Yu et al. 2024) 9,669 38.6375 12.7091 0.9371 39.4900 9.7035 0.9493 43.9150 2.9429 0.9704
VisionMamba[ICML 2024](Zhu et al. 2024) 27,880 38.7805 12.2990 0.9392 39.6339 9.3781 0.9506 43.9521 2.9103 0.9704
IRSRMamba[TGRS 2025 SOTA] (Huang et al. 2025) 26,462 39.3489 10.8767 0.9440 40.2302 8.3164 0.9548 44.5310 2.5537 0.9720
MambaIRv2 [CVPR 2025 SOTA] (Guo et al. 2025) 22,903 39.0421 11.5690 0.9423 39.9295 8.7817 0.9531 44.1442 2.7995 0.9711
GPSMamba (Ours) 36,942 39.3505 10.8608 0.9440 40.2418 8.2943 0.9548 44.5543 2.5451 0.9719
$\times 4$ EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 34.5219 30.1273 0.8548 35.1740 23.9917 0.8723 40.1190 6.8819 0.9482
ESRGAN[ECCVW 2018] (Wang et al. 2018) 16,661 33.6895 34.7337 0.8500 34.1650 28.9017 0.8679 37.9780 10.9641 0.9455
FSRCNN[ECCV 2016] (Dong et al. 2016) 475 33.8556 34.4909 0.8446 34.5272 27.4495 0.8636 38.7856 9.5482 0.9421
SRGAN[CVPR 2017] (Ledig et al. 2017) 1,370 34.5807 29.6927 0.8556 35.2076 23.7701 0.8728 40.1479 6.8162 0.9483
SwinIR[ICCV 2021] (Liang et al. 2021) 11,752 34.4321 30.6081 0.8537 35.0329 24.6490 0.8710 39.9062 7.1886 0.9479
SRCNN[T-PAMI 2015] (Dong et al. 2015) 57 33.6839 34.9181 0.8415 34.2348 28.6115 0.8568 38.0976 10.7588 0.9279
RCAN[ECCV 2018] (Zhang et al. 2018) 12,467 34.4280 30.8815 0.8528 35.0823 24.6507 0.8705 40.0805 6.9225 0.9484
PSRGAN[SPL 2021] (Huang et al. 2021) 2,414 34.4595 30.3760 0.8540 35.1023 24.3147 0.8715 39.9533 7.1274 0.9471
ShuffleMixer(tiny)[NIPS’22] (Sun et al. 2022)) 108 34.5440 29.9449 0.8550 35.1640 23.9705 0.8723 40.0756 6.9296 0.9478
ShuffleMixer (base)[NIPS’22] (Sun et al. 2022) 121 34.4507 30.6955 0.8538 35.0911 24.3745 0.8714 40.0120 7.0622 0.9477
HAT[CVPR 2023] (Chen et al. 2023) 20,624 34.4947 30.4086 0.8542 35.1239 24.3103 0.8713 40.0934 6.9078 0.9478
RGT[ICLR 2024] (Chen et al. 2024) 10,051 34.3826 31.0046 0.8535 35.0534 24.5924 0.8711 39.8420 7.3060 0.9472
MambaIR[ECCV 2024 SOTA] (Guo et al. 2024) 20,421 34.0267 32.9760 0.8510 34.5662 27.0850 0.8681 38.1878 10.8653 0.9404
ATD[CVPR 2024] (Zhang et al. 2024b) 753 34.6113 29.5294 0.8569 35.2347 23.6882 0.8737 40.2897 6.6001 0.9494
CATANet [CVPR 2025 SOTA] (Liu et al. 2025) 477 34.0613 32.7100 0.8447 34.8947 25.3673 0.8663 39.0302 8.8215 0.9405
MambaOut[CVPR 2025 SOTA](Yu et al. 2024) 9,669 34.4483 30.6792 0.8527 35.0456 24.7055 0.8698 40.1587 6.8076 0.9485
VisionMamba[ICML 2024](Zhu et al. 2024) 27,880 34.5941 29.5650 0.8564 35.2327 23.6467 0.8733 40.1705 6.7933 0.9484
IRSRMamba[TGRS 2025 SOTA] (Huang et al. 2025) 26,462 34.6755 29.0551 0.8577 35.3074 23.2857 0.8745 40.4052 6.4536 0.9497
MambaIRv2 [CVPR 2025 SOTA] (Guo et al. 2025) 22,903 29.5295 97.3011 0.8420 28.6532 130.8167 0.8577 27.2805 121.7986 0.9315
GPSMamba (Ours) 36,942 34.7421 28.7268 0.8587 35.4007 22.9636 0.8756 40.5475 6.2681 0.9503
TSAPC Loss: $\mathcal{L}_{T}$
In IRSR, conventional pixel-wise losses (e.g., L1 or L2) often fail to capture global structural coherence and can struggle to reconstruct faint but critical thermal signatures. This limitation is particularly pronounced for sequential models, such as Mamba, which benefit from global supervisory signals. To address this, we introduce the Thermal-Spectral Attention and Phase Consistency (TSAPC) Loss, a composite objective function designed to enhance both structural fidelity and the representation of salient thermal features in the reconstructed image $I_{SR}$.
Our proposed loss, denoted as $\mathcal{L}_{\text{T}}$, is a weighted sum of two frequency-domain components:
$\mathcal{L}_{\text{T}}=\lambda_{\text{phase}}\mathcal{L}_{\text{phase}}+\lambda_{\text{freq}}\mathcal{L}_{\text{freq}}$ (5)
where $\mathcal{L}_{\text{freq}}$ and $\mathcal{L}_{\text{phase}}$ are the two loss components, and $\lambda_{\text{freq}}$ and $\lambda_{\text{phase}}$ are their respective scalar weights. We now detail each component.
The first component, the Phase Consistency Loss ($\mathcal{L}_{\text{phase}}$), operates in the Fourier domain to enforce high-level structural correspondence between the $I_{SR}$ and $I_{HR}$. It is defined as the L1 distance between their phase angles:
$\mathcal{L}_{\text{phase}}=\left\|\text{angle}(F_{FFT}(I_{SR}))-\text{angle}(F_{FFT}(I_{HR}))\right\|_{1}$ (6)
where $F_{FFT}(\cdot)$ represents the 2D FFT and $\text{angle}(\cdot)$ extracts the phase component. Since image phase is known to carry critical information about structural layout, this loss term is vital for preserving sharp edges, object boundaries, and overall structural integrity in the reconstructed output.
The second component, the Thermal-Spectral Attention Loss ($\mathcal{L}_{\text{freq}}$), is engineered to focus the model’s learning on accurately reconstructing the spectral magnitudes within thermally significant regions. This is achieved by first generating a thermal attention mask, $M_{\text{thermal}}$, from $I_{HR}$. Specifically, $I_{HR}$ is passed through the feature extraction layers of a pre-trained VGG19 network ($\Phi_{\text{VGG}}$), and the resulting features are fed into a lightweight convolutional gating network $G_{\text{gate}}$ with a final sigmoid activation $\sigma(\cdot)$ to produce the mask:
$M_{\text{thermal}}=\sigma\left(G_{\text{gate}}\left(\Phi_{\text{VGG}}(I_{HR})\right)\right)$ (7)
This mask highlights regions of high thermal activity. After being interpolated to match the spatial dimensions of the images, $M_{\text{thermal}}$ is element-wise multiplied with both $I_{SR}$ and $I_{HR}$. The $\mathcal{L}_{\text{freq}}$ is then computed as the L1 distance between the magnitudes of the FFT of these masked images:
$\mathcal{L}_{\text{freq}}=\left\|\left|F_{FFT}(I_{SR}\odot M_{\text{thermal}})\right|-\left|F_{FFT}(I_{HR}\odot M_{\text{thermal}})\right|\right\|_{1}$ (8)
where $\odot$ denotes element-wise multiplication and $|\cdot|$ computes the magnitude. This semantically-guided spectral matching compels our model to prioritize the fidelity of frequency components originating from regions identified as thermally salient.
The synergy between $\mathcal{L}_{\text{phase}}$ and $\mathcal{L}_{\text{freq}}$ provides our GPSMamba architecture with both global structural constraints and feature-specific spectral guidance. This dual-component supervision is particularly beneficial for our Mamba-based architecture; while the inherent sequential processing excels at local dependencies, our proposed loss introduces crucial non-local, global supervisory signals. This helps the model transcend the limitations of its causal receptive field, encouraging it to learn representations coherent with the global context of the IR scene. Consequently, the TSAPC-Loss promotes the generation of images with superior detail recovery, enhanced contrast in thermally active regions, and more faithful rendering of complex thermal phenomena.
Experiments
In this section, we present a comprehensive evaluation of our proposed GPSMamba. We first detail the experimental setup, including datasets, evaluation metrics, and implementation specifics. We then provide extensive quantitative comparisons against SOTA methods, followed by qualitative visual analysis and in-depth ablation studies to validate our design choices.
Datasets. To ensure a rigorous and fair comparison with SOTA methods, we adopt the experimental protocol established by recent works like IRSRMamba. Our model was trained on the widely-used M3FD dataset(Liu et al. 2022b). For evaluation, we benchmark our GPSMamba on three standard infrared test sets: result-A(Liu 2018; Huang et al. 2021), result-C(Zhang et al. 2017; Huang et al. 2021), and CVC10(Campo et al. 2012). Following standard practice, the LR inputs were generated by applying bicubic downsampling to the HR images for scale factors of $\times 2$ and $\times 4$.
Evaluation Metrics. We assess performance using standard full-reference metrics: Peak Signal-to-Noise Ratio (PSNR/dB) and Structural Similarity Index (SSIM), computed on the Y channel of the YCbCr space. To provide a more holistic assessment of perceptual quality, which is not fully captured by pixel-wise measures, we supplement our analysis with four diverse no-reference (NR-IQA) metrics: the learning-based DBCNN(Zhang et al. 2020), the statistical NIQE(Mittal, Soundararajan, and Bovik 2012) and BRISQUE(Mittal, Moorthy, and Bovik 2012), and the vision-language based CLIP-IQA(Wang, Chan, and Loy 2023). Other Settings. All models were implemented in PyTorch and trained on an NVIDIA A6000 GPU. We optimized the network using the Adam optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 32.
(a) Error Distribution Across Models on result-A (Scale $\times 4$)
(b) Error Distribution Across Models on CVC10 (Scale $\times 4$)
Figure 3: Global error distribution analysis for various super-resolution models on the result-A and CVC10 datasets at scaling factor of $\times 4$. Each subfigure presents a comparative error breakdown across different ranges, quantifying the proportion of minimal, moderate, high, and severe reconstruction errors.
Experiment Results and Analysis
Quantitative Comparison.
Figure 4: Perceptual quality assessment of super-resolution models on result-A. Larger area indicates better perceptual performance. (Scale $\times 2$)
We conduct a comprehensive quantitative comparison against a suite of SOTA methods, including recent top-conference models like CATANet, MambaOut, and the current leading architecture, IRSRMamba. The full results for $\times 2$ and $\times 4$ super-resolution on three benchmark datasets are presented in Table 1. Our proposed model, GPSMamba, demonstrates exceptional performance. For the standard $\times 2$ SR task, it achieves either the best or second-best results across all datasets and metrics, establishing its competitive standing. For instance, on the result-C dataset, our method surpasses IRSRMamba in PSNR and matches its leading SSIM score. The superiority of our approach becomes more pronounced in the more challenging $\times 4$ SR task, where recovering high-frequency details is significantly more difficult. As shown in Table 1, our GPSMamba consistently sets a new SOTA, outperforming all competing methods on all metrics across all three datasets. On the result-C dataset, we achieve a PSNR of 35.4007 dB, surpassing the previous best (IRSRMamba) by a significant margin of nearly 0.1 dB. This marked improvement underscores our model’s enhanced capability to reconstruct faithful and structurally coherent images from highly degraded inputs. To dissect performance beyond aggregate metrics, we analyze the pixel-wise error distribution in Fig. 3. The visualization for the $\times 4$ task reveals our model’s advantage: it reconstructs 63% of pixels with minimal error ([0, 5)), surpassing all competitors (result-A (Scale $\times 4$)). Consequently, it produces the fewest pixels in the high-error categories ([10, $\infty$)). To complement our fidelity-based analysis, we assess perceptual quality using four non-reference metrics. As illustrated in Fig. 4, our method’s superiority is most evident on the challenging CLIP-IQA metric, where it achieves a leading score of 0.3043 and surpasses the current SOTA, IRSRMamba (0.2887). This advantage, which points to a more semantically coherent and realistic reconstruction, is supported by highly competitive performance across DBCNN, BRISQUE, and NIQE. The radar chart visually synthesizes these findings, showing our model’s performance envelope (blue) achieves the most favorable overall balance.
Figure 5: Visual comparison for $\times 4$ super-resolution on images fused2 (top) and result12 (bottom). Ours GPSMamba reconstructs significantly sharper details, a claim substantiated by the minimal reconstruction errors (darker blue) shown in the corresponding residual maps. Best view in zoom.
Table 2: Ablation studies validating our model design. We show the effectiveness of our core components (a) and justify our selection of key hyperparameters (b-d). Gray cells denote the final configuration. Metrics are PSNR/SSIM on result-A/result-C.
(a) (b) (c) (d)
PSNR SSIM Bach size PSNR SSIM Blocks PSNR SSIM $\lambda_{\text{freq}}$ $\lambda_{\text{phase}}$ PSNR SSIM
Mamba (Pure) 39.0346/39.9112 0.9417/0.9528 16 38.4182/39.1784 0.9417/0.9440 4 38.9902/39.9084 0.9428/0.9536 0.2 0.8 39.3157/40.2283 0.9440/0.9548
$\mathbf{w}\text{ASF-SSM}$ 39.1472/40.0727 0.9429/0.9716 32 39.1681/40.1465 0.9520/0.9547 6 39.0614/40.0688 0.9433/0.9540 0.5 0.3 39.3346/40.2236 0.9439/0.9547
$\mathbf{w}\text{ASF-SSM}\&\mathscr{L}_{T}$ 39.3142/40.2442 0.9440/0.9548 64 39.0520/40.0754 0.9434/0.9542 8 39.1030/40.0934 0.9442/0.9549 0.8 0.2 39.3410/40.2216 0.9439/0.9547
Figure 6: Impact of the Axial Shifted Fusion State Space Module (ASF-SSM) on the Effective Receptive Field (ERF).
Qualitative Results. As in Fig.5, qualitative comparison for $\times 4$ super-resolution on the result-A/result-C. Our model, GPSMamba, excels in restoring sharp textures and object contours where other methods fail. In the fused2 scene (top), it renders crisp, legible text on the sign, a detail completely lost to blurring in competing results. Similarly, in result12 (bottom), it preserves the pedestrian’s sharp silhouette. This superior fidelity is further substantiated by the residual maps (right), where our method exhibits minimal reconstruction error (darker blue) in these critical high-frequency regions, confirming a more accurate reconstruction.
Ablation Experiments. As shown in Table.2(a), we begin with a Mamba (Pure) baseline. Integrating our proposed ASF-SSM yields a significant performance boost, increasing PSNR from 39.03 to 39.14 and SSIM from 0.9528 to 0.9716 on result-A/result-C. This demonstrates the module’s efficacy in capturing critical image features. By further incorporating our total loss function ($\mathscr{L}_{T}$), which includes spectral constraints, the performance is elevated to its peak (39.31/40.24 PSNR), confirming that each component contributes synergistically to the final result. We also analyze the impact of key hyperparameters. As detailed in Table.2(b), we test batch sizes of 16, 32, 64. The model achieves optimal performance with a batch size of 32, which we adopt for all experiments. Table.2(c) shows the effect of varying the number of our core blocks. Performance consistently improves as the model deepens from 4 to 8 blocks, with 8 blocks delivering the best results. Finally, we study the balance between the frequency ($\lambda_{\text{freq}}$) and phase ($\lambda_{\text{phase}}$) loss terms in Table.2(d). The results indicate that placing a higher emphasis on the frequency component is beneficial, with the optimal configuration being $\lambda_{\text{freq}}=0.8$ and $\lambda_{\text{phase}}=0.2$.
Effect of ASF-SSM and TSAPC Loss. We visually validate our core components through an ablation study. As shown in Fig.6, our ASF-SSM transforms the model’s Effective Receptive Field(Ding et al. 2022; Li et al. 2023) from a compact, localized region (w/o ASF-SSM) to an expanded and anisotropic one (w ASF-SSM), confirming its capacity for capturing long-range dependencies. Concurrently, the frequency analysis in Fig.7 demonstrates that our TSAPC loss is critical for structural fidelity. Without it (w/o $\mathcal{L}_{\text{T}}$), the reconstructed phase spectrum exhibits severe artifacts. By enforcing spectral consistency, our method (w $\mathcal{L}_{\text{T}}$) produces a phase spectrum that faithfully aligns with the ground truth, leading to perceptually superior reconstructions.
Figure 7: Visual comparison in spatial and frequency domains. From top to bottom: Ground Truth, baseline model (w/o $\mathcal{L}_{\text{T}}$), and our full model (w $\mathcal{L}_{\text{T}}$).
Conclusion
We proposed GPSMamba, an IRSR architecture designed to overcome the global context fragmentation inherent in the 1D causal scanning of SSMs. GPSMamba introduces two core innovations: the ASF-SSM, which injects a fused semantic systematic prompt to guide reconstruction with non-local context, and the TSAPC Loss ($\mathcal{L}_{T}$), which provides explicit, non-causal supervision to enforce structural and spectral fidelity. Our proposed GPSMamba also achieves state-of-the-art performance, demonstrating that the causal modeling limitations of SSMs can be effectively overcome for visual representation learning.